# Phase space maximal entropy random walk: Langevin-like ensemble of physical trajectories

### v1

@@ III-A MERW continuous limit in phase space
In one step $\tilde{x}\to\tilde{x}+\tilde{v}$, and there is a random transition of $\tilde{v}$, we can assume at most ±$\pm 1$. To include damping in future, we could add velocity reduction in below step. In 1D the MERW eigenequation after neglecting higher order terms becomes:
$\lambda\psi_{\tilde{x}-\tilde{v},\tilde{v}}=(M\psi)_{\tilde{x}-\tilde{v},\tilde{v}}=\sum_{s\in\{-1,0,1\}}e^{-\epsilon\beta V_{\tilde{x},\tilde{v}+s}}\psi_{\tilde{x},\tilde{v}+s}\approx$ (12)
$\approx\psi_{\tilde{x},\tilde{v}-1}+\psi_{\tilde{x},\tilde{v}}+\psi_{\tilde{x},\tilde{v}+1}-3\beta\epsilon V_{\tilde{x},\tilde{v}}\psi_{\tilde{x},\tilde{v}}$
Approximating using derivative $\nabla_{x}\equiv\partial/\partial x$:
$\psi_{\tilde{x}-\tilde{v},\tilde{v}}=\psi(\delta(\tilde{x}-\tilde{v}),v)=\psi(x-\epsilon v,v)\approx(\psi-\epsilon v\nabla_{x}\psi)_{\tilde{x},\tilde{v}}$
and substituting it to eigenequation (12) we get:
$\lambda(\psi-\epsilon v\nabla_{x}\psi)_{\tilde{x},\tilde{v}}\approx\psi_{\tilde{x},\tilde{v}-1}+\psi_{\tilde{x},\tilde{v}}+\psi_{\tilde{x},\tilde{v}+1}-3\beta\epsilon V_{\tilde{x},\tilde{v}}\psi_{\tilde{x},\tilde{v}}$
@@
$\frac{3-\lambda}{3\beta(\delta/\epsilon)^{2}}\psi_{\tilde{x},\tilde{v}}+v\frac{\lambda}{3\beta}\frac{\epsilon}{(\delta/\epsilon)^{2}}(\nabla_{x}\psi)_{\tilde{x},\tilde{v}}\approx$
$\approx-\frac{1}{3\beta}\frac{\psi_{\tilde{x},\tilde{v}-1}-2\psi_{\tilde{x},\tilde{v}}+\psi_{\tilde{x},\tilde{v}+1}}{(\delta/\epsilon)^{2}}+\frac{\epsilon}{(\delta/\epsilon)^{2}}V_{\tilde{x},\tilde{v}}\psi_{\tilde{x},\tilde{v}}$
@@
For $\nabla_{x}$ gradient term we can choose $(\lambda\to 3)$:
$\frac{\lambda}{3\beta}\frac{\epsilon}{(\delta/\epsilon)^{2}}\to\xi\qquad\textrm{for example}\quad\epsilon=\sqrt[3]{\beta\xi\,\delta^{2}}$ (13)
@@
$E\psi=-C\Delta_{v}\psi-\xi v\cdot\nabla_{x}\psi+(V+m\|v\|^{2}/2)\psi$ (14)
@@
To find the stationary probability distribution $\rho(x,v)$, we need to solve eigenequation (14) minimizing energy $E$ (higher energy contributions have exponential decay), then $\rho(x,v)\propto\psi(x,v)^{2}$.
@@
While for QM/MERW there are well known analytical solutions of such stationary Schrödinger equation e.g. for harmonic potential, infinite well, point $-1/r$ potential, their phase space version are more complicated and Wolfram Mathematica DEigensystem was not able to find analytical solutions, hence for their example solution in Fig. 1 there was used numerical NDEigensystem, also required for more complex settings.
@@
Function NDEigensystem uses finite element method, and already had various issues in the discussed basic cases - resolved e.g. by manually specifying Dirchlet conditions for infinite well, or for point potential: solving for $\psi(x,v)\sim\exp(-v^{2}-x)\phi(x,v)$ type substitution explicitly assuming asymptotic behavior. Alternative approach is approximation by solving in a finite orthogonal basis of functions (removing higher modes), especially Gaussian distribution times Hermite polynomial for velocities (ground state already gets such higher modes also for velocity). Numerical methods for psMERW will require further work, especially to include interaction between multiple particles.
IV Conclusions and further work
@@
While popularly considered path ensembles usually use infinite kinetic energy paths, it is mathematically more difficult but doable to use ensembles of more physical paths instead by going to phase space. This article discusses some basics with included spatial potential to bring attention, hopefully leading to some attempts to experimentally determine the most appropriate ones for various physical settings.
@@
This is initial article opening various directions for further work, e.g.:
@@
  • Search for possibilities of experimental determination which of the discussed approaches is the most appropriate, especially using stationary probability distributions of positions, maybe also velocities e.g. from redshifts - in various settings from microscopic e.g. molecules, solitons to astronomical e.g. dust halos.
@@
  • Mathematical improvements, especially of numerical approaches, also search for analytical formulas for basic cases, maybe include damping like in Langevin equation, larger velocity changes through collisions, include relativistic corrections, etc.
  • While for simplicity we have discussed only stationary situation, it seems worth to also consider dynamics - both straightforward evolution in fixed potential like for electron conductance [5], but also much more difficult case of varying potential - with basics discussed in [9].
  • While the discussed analysis was for a single walker, in practice we usually have multiple interacting - what should be finally included. One way is through mean-field treatment assuming the remaining have the same distribution, used e.g. for MERW conductance model [5] to include potential contributions from the remaining electrons. For psMERW it could be useful e.g. for astronomical dust halos to include their gravitational self-interaction.

### v4

@@ III-B1 Derivation
In one step $\tilde{x}\to\tilde{x}+\tilde{v}$, and there is a random velocity change $\tilde{v}\to\tilde{v}+s$, we can assume $s\in\{-1,0,1\}$. To include damping in future, we could add velocity reduction in below step. In 1D the MERW eigenequation after neglecting $O(\epsilon^{2})$ terms becomes:
$\lambda\psi_{\tilde{x},\tilde{v}}=(M\psi)_{\tilde{x},\tilde{v}}=\sum_{s\in\{-1,0,1\}}e^{-\epsilon\beta\bar{V}_{\tilde{x}+\tilde{v},\tilde{v}+s}}\psi_{\tilde{x}+\tilde{v},\tilde{v}+s}\approx$ (13)
$\approx\sum_{s\in\{-1,0,1\}}(1-\epsilon\beta\bar{V}_{\tilde{x},\tilde{v}})(\psi_{\tilde{x},\tilde{v}+s}+\epsilon v(\nabla_{x}\psi)_{\tilde{x},\tilde{v}})$
thanks to approximation with derivative $\nabla_{x}\equiv\partial/\partial x$:
$\psi_{\tilde{x}+\tilde{v},\tilde{v}}=\psi(\delta(\tilde{x}+\tilde{v}),v)=\psi(x+\epsilon v,v)\approx\psi_{\tilde{x},\tilde{v}}+\epsilon v(\nabla_{x}\psi)_{\tilde{x},\tilde{v}}$
Neglecting $O(\epsilon^{2})$ terms eigenequation (13) becomes:
$\lambda\psi_{\tilde{x},\tilde{v}}\approx\psi_{\tilde{x},\tilde{v}-1}+\psi_{\tilde{x},\tilde{v}}+\psi_{\tilde{x},\tilde{v}+1}+3\epsilon v(\nabla_{x}\psi)_{\tilde{x},\tilde{v}}-3\beta\epsilon\bar{V}_{\tilde{x},\tilde{v}}\psi_{\tilde{x},\tilde{v}}$
@@
$\frac{3-\lambda}{3\beta(\delta/\epsilon)^{2}}\psi_{\tilde{x},\tilde{v}}\approx-\frac{1}{3\beta}\frac{\psi_{\tilde{x},\tilde{v}-1}-2\psi_{\tilde{x},\tilde{v}}+\psi_{\tilde{x},\tilde{v}+1}}{(\delta/\epsilon)^{2}}-$
$-v\frac{1}{\beta}\frac{\epsilon}{(\delta/\epsilon)^{2}}(\nabla_{x}\psi)_{\tilde{x},\tilde{v}}+\frac{\epsilon}{(\delta/\epsilon)^{2}}\bar{V}_{\tilde{x},\tilde{v}}\psi_{\tilde{x},\tilde{v}}$
@@
For $\nabla_{x}$ gradient term we can choose:
$\frac{1}{\beta}\frac{\epsilon}{(\delta/\epsilon)^{2}}\to\xi\qquad\textrm{for example}\quad\epsilon=\sqrt[3]{\beta\xi\,\delta^{2}}$ (14)
@@
$E\psi=-C\Delta_{v}\psi-\xi v\cdot\nabla_{x}\psi+\left(V+\frac{1}{2}m\|v\|^{2}\right)\psi$ (15)
@@
To find the stationary probability distribution $\rho(x,v)$, we need to solve eigenequation (15) minimizing energy $E$, what corresponds to $|\lambda|$ maximization. Higher energy contributions have exponential decay here. This time $M$ is not symmetric as $M^{T}$ time symmetry inverses velocity, so we should find left/right eigenfunction. Fortunately, as in Fig. 6, they differ by just change of sign of velocity $\phi(x,v)=\psi(x,-v)$, getting Born rule:
$\rho(x,v)\propto\psi(x,v)\ \psi(x,-v)$ (16)
III-B2 Comparison with Feynman path ensemble
Similar $t\to it$ ”Wick rotated” phase space Schrödinger equation was earlier derived by Bouchaud [11] for QM, Feynman path ensembles:
$i\hbar\frac{\partial\Psi(x,q)}{\partial t}=\frac{\hbar}{2\tau_{c}}[\Delta_{q}\Psi(x,q)-q^{2}\Psi(x,q)]+$
$\qquad+i\sqrt{\frac{\hbar^{3}}{m\tau_{c}}}q\nabla_{x}\Psi(x,q)+V(x)\Psi(x,q)$ (17)
for $\tau_{c}$ momentum correlation time: as in Ornstein-Uhlenbeck process assuming $\langle p(s)p(u)\rangle=\frac{m\hbar}{2\tau_{c}}\exp(-|s-u|/\tau_{c})$, and rescaled momentum $p=\sqrt{m\hbar/\tau_{c}}q$. Comparing it with (15) suggests parameter choice for quantum scale applications.
Beside $t\to it$ ”Wick rotation to imaginary time”, there are some subtle differences between Boltzmann and Feynman path ensembles. Hamiltonian corresponds to minus $M$ matrix, energy $E$ corresponds to minus $\lambda$ discussed here maximization of $\lambda$ corresponds to minimization of energy for the ground state.
In Boltzmann path ensemble excited states vanish exponentially: $M\psi^{k}=\lambda_{k}\psi^{k}$ for $|\lambda_{0}|\geq|\lambda_{1}|\geq\ldots$, then after $t$ steps $\sum_{k}a_{k}\psi^{k}\to\sum_{k}(\lambda_{k}/\lambda_{0})^{t}a_{k}\psi^{k}$ (+normalization). In contrast, in QM/Feynman path ensembles $\psi^{k}$ excited states are stable, rotate in complex plane. Excited e.g. atoms are believed to deexcite through interaction with environment.
Transposition $M^{T}$ corresponds to time reversal. As in Fig. 6, for non-symmetric $M$ ensembles of future and past half-paths lead to different $\phi,\psi$ which are real for Boltzmann ensembles. In contrast, (17) Hamiltonian is self-adjoint: invariant under time symmetry in conjugation, containing such two asymmetric real eigenfunction in real/imaginary parts of one complex wavefunction.
III-B3 Calculation remarks
@@
While for QM/MERW there are well known analytical solutions of such stationary Schrödinger equation e.g. for harmonic potential, infinite well, point $V\propto-1/r$ potential, their phase space version are more complicated and Wolfram Mathematica DEigensystem was not able to find analytical solutions, hence for their example solution in Fig. 1 there was used numerical NDEigensystem, also required for more complex settings. It uses finite element method, and already had various issues in the discussed basic cases - resolved e.g. by manually specifying Dirchlet conditions for infinite well, or for point potential: solving for $\psi(x,v)\sim\exp(-v^{2}-x)\phi(x,v)$ type substitution assuming asymptotic behavior.
@@
Alternative approaches are discretization like in Fig. 2, or approximation by solving in a finite orthogonal basis of functions (removing higher modes), especially Gaussian distribution times Hermite polynomial for velocities (ground state already gets such higher modes), further developed in [11]. Numerical methods for psMERW will require further work, especially to include interaction between multiple particles.
IV Boltzmann ensembles of Lévy flights
Laplacians in the discussed derivations came from infinitesimal behavior of $x\to x\pm 1$ (MERW) or $v\to v\pm 1$ (psMERW) random steps. As these random steps have finite variance, the central limit theorem says they lead to Gaussian distribution, which can be deformed here through Boltzmann path ensemble.
Without the assumption of finite variance, there is generalized central limit theorem [17], which in infinitesimal limit replaces Laplacian with its Riesz fractional derivative generalization, leading to Lévy flights [18] with steps of infinite variance.
[…]
$\varphi_{\alpha s}(k)=\exp\left(it\delta-\|sk\|^{\alpha}\right)$ (18)
generally also with asymmetry term, usually neglected in Lévy flights (maybe it should be considered?). For $\alpha=2$ it is just Gaussian distribution, for $\alpha=1$ it is Cauchy distribution. For $\alpha<2$ it has infinite variance, and $\rho(x)\sim|x|^{-1-\alpha}$ tails.
While $x\to x\pm 1$ has lead to Laplacian in infinitesimal limit, in Lévy flights we replace such finite variance term with of $\exp\left(-\|k\|^{\alpha}\right)$ characteristic function - leading to fractional Riesz derivative, Laplacian, in practice defined for characteristic function:
$F_{k}\left[(-\Delta)^{\alpha/2}\varphi\right]=\|k\|^{\alpha}F_{k}[\varphi]$ (19)
where $\|\cdot\|\equiv\|\cdot\|_{2}$ is Euclidean norm - it is crucial to use spherically symmetric. Using coordinate-wise $|k|^{\alpha}$ instead, would lead to dominant steps in canonical directions.
For MERW as Boltzmann path ensemble, to use ensemble of Lévy trajectories of infinite step variance, we need to replace Laplacian in Schrödinger equation wits such fractional one $(-\Delta_{x})^{\alpha/2}$. Such fractional Schrödinger equation is considered in literature [19], also recently getting experimental realization [20]. For example in 1D infinite potential well, in MERW we get $EF_{k}[\psi]=|k|^{\alpha}F_{k}[\psi]$ type eigenequation, due to Dirchlet conditions having $\sin$-like solutions - going to Boltzmann ensemble of Levy trajectories does not change stationary probability distribution, only energy levels.
For psMERW analogously we can use Lévy flights of velocities by replacing Laplacian in (15) with fractional $(-\Delta_{v})^{\alpha/2}$, getting phase space fractional Schrödinger equation. It would allow to model ensembles of smooth paths with changes of velocity having infinite variance, for example corresponding to large velocity changes during collisions in e.g. dust. Applying Fourier transform to velocity $\tilde{\psi}\equiv\tilde{\psi}(x,k)=\int e^{ikv}\psi(x,v)$ it becomes:
[…]
$E\tilde{\psi}=(C|k|^{\alpha}+V)\tilde{\psi}-\frac{m}{2}\partial_{kk}\tilde{\psi}+i\xi\partial_{xk}\tilde{\psi}$ (20)
Figure 7: Tunneling investigation setting (top), Mathematica source (center), and simulation results (bottom) of tunneling probability dependence from barrier width $w$ and initial velocity $v$ (and number of steps in bottom-right). The (discrete) setting is potential barrier of varying width, with symmetric two zero potential regions of the same fixed width on both sides (ending with infinite potential walls). We start with probability distribution concentrated in a single point just before the barrier (green start), and chosen initial velocity (working in phase space). Then we evolve this distributions using the found psMERW stochastic propagator, zeroing all density which visited the two absorption regions (absorb and finish), earlier storing probabilities which ended in the finish region. There are shown plots for fixed initial velocity (upper-right) and on its left using rough approximation with guessed $\sim 1-\textrm{sech}(v^{2}/w)$ formula. In logarithmic scale we can observe constant asymptotic direction: exponent of transition probability reduction with barrier width $w$. Bottom-left plot fixes barrier width instead, showing transition probability approaching 1 for high initial velocities. Similar plots were found in 2009 in walking droplet tunneling experiment [21, 22, 23]. Bottom-right plot fixes barrier width $5002w=$ 500, showing probability of reaching the finish in a given number of steps - e.g. allowing to estimate mean velocity: lowered for fast initial velocity, increased for low initial.
V Tunneling simulations with initial velocity
As in 2, we can simulate tunneling with the discussed path ensembles - for example using infinite potential well with additional barrier inside, e.g. rectangular assumed here.
[…]
In literature, looking similar dependencies were found especially in 2009 tunneling experiment for walking droplets [21, 22, 23]. They complement classical behavior with history dependence through field carrying (pilot-like) waves from the previous bounces - high complexity of this dependence leads to practically random transition probabilities. Not knowing details of such hidden field, from statistical physics perspective it seems natural to assume effective description by Boltzmann path ensembles - requiring smooth paths: phase space variants.
This behavior can be approximated by ordinary differential equations - considered for tunneling in [24]. However, it leads to much more complicated behavior, currently not getting the required exponential decrease of transition probabilities with barrier width.
The details of such initial velocity dependence of tunneling probability were not found in literature - requiring further investigation, maybe deriving from the discussed smooth path ensembles, phase space Schrödinger equation, or trying to fit the numerical simulations - which is close to $\approx 1-\textrm{sech}(av^{2}/w)$ type dependence for some $a$ parameter.
VI Conclusions and further work
@@
While popularly considered path ensembles usually use infinite kinetic energy paths, it is mathematically more difficult but doable to use ensembles of more physical paths instead by going to phase space, however, it leads to slightly different predictions. This article discusses some basics with included spatial potential to bring attention, hopefully leading to some attempts to experimentally determine the most appropriate ones for various physical settings.
@@
This is early article opening various directions for further work, e.g.:
@@
  • Search for possibilities of experimental determination which of the discussed approaches is the most appropriate for various scenarios, especially using stationary probability distributions of positions, maybe also velocities e.g. from redshifts, or in dynamical settings like tunneling - in various situations from microscopic e.g. neutrons [7], molecules, solitons [8], to astronomical e.g. dust halos.
@@
  • Mathematical improvements, especially of numerical approaches, search for analytical formulas for basic cases, maybe include damping like in Langevin equation, larger velocity changes through collisions, relativistic corrections, magnetic field through vector potential, etc.
  • While for simplicity we have discussed only stationary situation, it is worth to also consider dynamics - both evolution in fixed potential like for electron conductance [6] or tunneling in Fig. 2, but also much more difficult case of varying potential - with basics discussed in [15].
  • While the discussed analysis was for a single walker, in practice we usually have multiple interacting - what should be finally included. One way is through mean-field treatment assuming the remaining have the same distribution, used e.g. for MERW electron conductance model [6] to include potential contributions from the remaining electrons. For psMERW it could be useful e.g. for astronomical dust halos to include their gravitational self-interaction.
  • While we have focused on ensembles of full paths, maybe it is also valuable to consider ensembles of unidirectional paths, which have asymmetric velocity distributions like in Fig. 6.
@@
  • Investigate tunneling probability dependence - not only from barrier widths, but also initial velocity.